# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

### v1

@@ E.4 Comparing all variants of CLEAR
Appendix F Case Study: Housing Market Price Prediction and changing the number of predictors
@@
Table 1 and the discussion in Section 4.3 summarize the results. In the reduced 2-variable setting, CLEAR selected $\lambda=0.6$, assigning more weight to aleatoric uncertainty. In contrast, in the full-feature setting, $\lambda=14.5$, emphasizing epistemic uncertainty. The corresponding calibrated epistemic-to-aleatoric ratios were computed as $\frac{1}{n}\sum_{i=1}^{n}\frac{\gamma_{1}\times\text{epistemic}(x_{i})}{\gamma_{2}\times\text{aleatoric}(x_{i})}$ in (1), reflecting the final width contributions from each component. They were 0.03 and 7.72, respectively. This ratio quantifies how much of the interval width is attributed to epistemic versus aleatoric uncertainty after calibration. The results demonstrate CLEAR’s ability to adaptively reweight the two uncertainty sources in response to the underlying information regime, producing intervals that are both sharp and well-calibrated across heterogeneous settings.
Appendix G Theory: coverage guarantees for CLEAR
G.1 Finite-sample marginal coverage for CLEAR
  1. 1.
  2. 2.
However, our empirical results in Tables 6, 11 and 16 show that for reasonably sized datasets, this theoretical discrepancy does not visibly impact our practical marginal coverage. From a theoretical perspective, CLEAR still achieves asymptotic marginal coverage (as a consequence of asymptotic conditional coverage, see Lemma 2.1).
Definition G.1 (Conformalized CLEAR).
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Lemma G.2.
$\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}}))\geq 1-\alpha.$
Proof.
However, in practice, we recommend using our default implementation of CLEAR, where the validation data is reused for calibration ($\mathcal{D}_{\text{cal}}=\mathcal{D}_{\text{val}}$). This improves data efficiency while still achieving strong approximate marginal coverage in our experiments (see Tables 6, 11 and 16). Avoiding a split allows more data for both validation and calibration, which improves model selection and stabilizes marginal and conditional coverage. Even when marginal coverage $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}}))\geq 1-\alpha$ is satisfied, the actual coverage $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{cal}})$ can be significantly lower than the target coverage $1-\alpha$ for a fixed calibration dataset $\mathcal{D}_{\text{cal}}$, due to the inherent variation of $\mathcal{D}_{\text{cal}}$. Only with sufficiently large calibration datasets $\mathcal{D}_{\text{cal}}$ can we expect $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{cal}})$ to be reliably close to $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}}))$. [^4] Conformal theory provides theoretical guarantees for $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}}))\geq 1-\alpha$ and $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}})\geq 1-\alpha$ under the standard exchangeability assumption. However, it does not provide finite-sample guarantees for $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{cal}})$ and $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{cal}})$ for a fixed calibration dataset, even under the standard exchangeability assumption.
G.2 Asymptotic conditional coverage for CLEAR — discussion of the assumptions in Lemma 2.1
  1. 1.
  2. 2.
  3. 3.
Appendix H On the Role of Relative and Absolute Uncertainty in Coverage Guarantees
@@ H.1 Metrics and their focus
  • NCIW evaluates the ranking of uncertainty—whether the method assigns wider intervals to more uncertain points—but is invariant to the overall scale. It reflects relative uncertainty without penalizing miscalibration. Minimum Negative Log-Likelihood (NLLmin) [28] similarly measures relative uncertainty, focusing on whether the method correctly ranks more vs. less uncertain inputs, independent of the predicted scale.
  • Quantile Loss, CRPS, and AISL combine both relative and absolute components. These metrics penalize both poor ranking and miscalibration, rewarding methods that adapt well to heteroskedasticity while maintaining valid coverage.
  • PICP and NIW each provide only partial information: PICP measures calibration but is blind to adaptivity, while NIW captures the average scale of uncertainty. However, optimizing both jointly implicitly requires the method to rank uncertainty well and calibrate it accurately.
@@
H.2 Applications
Understanding whether a method captures relative or absolute uncertainty has practical implications across a range of applications:
In active learning, the primary objective is to identify inputs $x$ for which the model is most uncertain, guiding efficient data acquisition. Here, only the ranking of uncertainty matters—selecting the point with the highest epistemic uncertainty. Methods that preserve good relative uncertainty, even if miscalibrated, often suffice.
In Bayesian optimization, many acquisition functions (such as upper confidence bound or entropy search) depend more on relative than absolute uncertainty [107, 108]. Using upper bounds of the form $\hat{f}(x)+c$ with constant $c$ across all $x$ does not improve over exploiting $\hat{f}(x)$ alone, highlighting the centrality of uncertainty ranking over calibration in this setting.
In human-in-the-loop automation, relative uncertainty can guide prioritization—for instance, flagging uncertain cases for expert review. While calibrated intervals may not always be necessary, correct ordering of confidence can improve decision efficiency and safety.
H.3 Methods
@@
H.4 Calibration Techniques
The absolute scale of uncertainty is critical in applications where the expected risk or cost over a population matters. For example, in probabilistic risk management, climate forecasting, or decision-making under uncertainty, one needs valid estimates of average confidence. However, as emphasized by pathological cases that achieve perfect PICP with no adaptivity, relative uncertainty remains essential for practical utility. Several approaches have been developed to adjust the scale of uncertainty estimates while preserving different structures:
@@ H.4 Calibration Techniques
  • Isotonic calibration [3] applies a nonparametric monotonic transformation that preserves the ranking of uncertainty, suitable when only order is trusted.
@@
H.5 Achieving Conditional Coverage
@@
Even when conditional coverage is not required, improving relative uncertainty tends to reduce average interval width and improve marginal calibration. Thus, adaptivity and calibration are not mutually exclusive but can reinforce one another.
CLEAR is explicitly designed to target both forms of uncertainty—modeling and calibrating relative and absolute epistemic, aleatoric, and total predictive uncertainty. As such, it provides a flexible and principled framework for applications demanding both adaptivity and reliability.

### v2

@@ G.3 Variant (c) Conformalized
Appendix H Case Study: House Price Prediction with Varying Number of Predictors
@@
Table 2 and the discussion in Section 4.3 summarize the results. In the reduced 2-variable setting, CLEAR selected $\lambda=0.6$, assigning more weight to aleatoric uncertainty. In contrast, in the full-feature setting, $\lambda=14.5$, emphasizing epistemic uncertainty. The corresponding calibrated epistemic-to-aleatoric ratios were computed as $\frac{1}{n}\sum_{i=1}^{n}\frac{\gamma_{1}\times\text{epistemic}(x_{i})}{\gamma_{2}\times\text{aleatoric}(x_{i})}$ in (1), reflecting the final width contributions from each component. They were 0.03 and 7.72, respectively. This ratio quantifies how much of the interval width is attributed to epistemic versus aleatoric uncertainty after calibration. The results demonstrate CLEAR’s ability to adaptively re-weight the two uncertainty sources in response to the underlying information regime, producing intervals that are both sharp and well-calibrated across heterogeneous settings.
Appendix I On the Role of Relative and Absolute Uncertainty in Coverage Guarantees
@@ I.1 Metrics: Relevance
  • Quantile Loss, CRPS, and AISL combine both relative and absolute components incentivizing conditional coverage. These metrics penalize both poor ranking and miscalibration, rewarding methods that adapt well to heteroskedasticity.
  • NCIW is invariant to the overall scale of the uncertainty but evaluates the ranking—whether a method assigns wider intervals to more uncertain points. In practice, a low NCIW often correlates with good relative uncertainty. However, a minimal NCIW theoretically encourages suboptimal conditional coverage, as it can lead to under-coverage in high-uncertainty regions and over-coverage in low-uncertainty regions. Consequently, good conditional coverage typically requires a slightly higher NCIW than the minimum. Similar to NCIW, Minimum Negative Log-Likelihood (NLLmin) [27] also assesses relative uncertainty by focusing on whether a method correctly ranks more versus less uncertain inputs, independent of the predicted scale.
  • PICP and NIW each provide only partial information: PICP measures calibration but is blind to adaptivity, while NIW captures the average scale of uncertainty.
@@
I.2 Applications
Understanding whether a method captures relative or absolute uncertainty has practical implications across a range of applications. In active learning, the primary objective is to identify inputs $x$ for which the model is most uncertain, guiding efficient data acquisition. Here, only the ranking of uncertainty matters—selecting the point with the highest epistemic uncertainty. Methods that preserve good relative uncertainty, even if miscalibrated, often suffice. In Bayesian optimization, many acquisition functions (such as upper confidence bound or entropy search) depend more on relative than absolute uncertainty [114, 115]. Using upper bounds of the form $\hat{f}(x)+c$ with constant $c$ across all $x$ does not improve over exploiting $\hat{f}(x)$ alone, highlighting the centrality of uncertainty ranking over calibration in this setting. In human-in-the-loop automation, relative uncertainty can guide prioritization—for instance, flagging uncertain cases for expert review. While calibrated intervals may not always be necessary, correct ordering of confidence can improve decision efficiency and safety.
I.3 Methods
@@
I.4 Calibration Techniques
The absolute scale of uncertainty is critical in applications where the expected risk or cost over a population matters (e.g., in probabilistic risk management, climate forecasting, or decision-making under uncertainty). However, as emphasized by pathological cases that achieve perfect PICP with no adaptivity, relative uncertainty remains essential for practicality. Several approaches can be taken to adjust the scale of uncertainty estimates while preserving different structures:
@@ I.4 Calibration Techniques
  • Isotonic calibration applies a nonparametric monotonic transformation that preserves the ranking of uncertainty, suitable when only order is trusted. E.g., [3] uses isotonic calibration for distributional uncertainty quantification.
@@
I.5 Achieving Conditional Coverage
@@
Even when conditional coverage is not required, improving relative uncertainty tends to reduce average interval width and improve marginal calibration under distribution shifts. Thus, adaptivity and calibration are not mutually exclusive but can reinforce one another. CLEAR is explicitly designed to target both forms of uncertainty—modeling and calibrating relative and absolute epistemic, aleatoric, and total predictive uncertainty. As such, it provides a flexible and principled framework for applications demanding both adaptivity and reliability.